Amy is 6 years older than Bill. 2 years ago Amy was 3 times as old as Bill. How old will Amy be in 4 years?
Answer by addingup(3677)

A = B+6
A-2 = 3B
B+6-2 = 3B
B+4 = 3B
4 = 2B
2B = 4
B = 2
A = 2+6 = 8 and in 4 years she will be 8+4 = 12